find the equation of the line, given the following information
Passes through (1,4) and is parallel to y=5x + 2
1
Expert's answer
2020-06-09T17:25:03-0400

The given line is,

y = 5x + 2

And the point through which the line parallel to the given line passes is, (1,4).


The slope of y = 5x + 2 is, m = 5.

The line parallel to the given line will also have the same slope.


Using point slope equation of line, we get,


(y - y1) = 5(x - x1)


Substituting the point (1,4), we get,


y - 4 = 5(x - 1)

y - 4 = 5x - 5

i.e., y = 5x - 5 + 4

"\therefore" y = 5x - 1


The line parallel to the line y = 5x + 2 that passes through the point (1,4) is, y = 5x - 1.